OpenStudy (anonymous):

What are the factors of x2 + 10x + 24? (x − 12)(x − 2) (x + 12)(x + 2) (x + 4)(x + 6) (x − 4)(x − 6)

OpenStudy (theviper):

\[=x^2+4x+6x+24\]\[=x(x+4)+6(x+4)\]\[=(x+6)(x+4)\]

OpenStudy (theviper):

So, C.

OpenStudy (anonymous):

typo error, (x+4)(x+6)

OpenStudy (theviper):

ha sahi kaha

OpenStudy (theviper):

yes

OpenStudy (theviper):

(x+4)(x+6)